OpenStudy (tylermckinney16):

Help please :) Prove or disprove that the quadrilateral defined by the points A(6,2), B(3,4), C(1,1), D(4,-1) is a rectangle.

OpenStudy (tylermckinney16):

@Directrix

OpenStudy (tylermckinney16):

Slope(m)=-0.667 This is the answer to the one u told me to find using the calculator.

OpenStudy (tylermckinney16):

@Directrix

Directrix (directrix):

Consider that to be -2/3 as the slope for segment AB.

Directrix (directrix):

What is the slope for segment BC? B(3,4), C(1,1) http://calculator.tutorvista.com/slope-calculator.html

OpenStudy (tylermckinney16):

Slope(m)= 1.5

Directrix (directrix):

Segment AB has slope -2/3. Segment BC has slope 3/2. -2/3 * 3/2 = -1. If the product of the slopes of two lines is -1, then the lines are perpendicular. So, Segment AB is perpendicular to segment BC.

Directrix (directrix):

What is the slope for segment CD? C(1,1), D(4,-1)

OpenStudy (tylermckinney16):

Slope(m)=-0.667

Directrix (directrix):

Slope of CD = =2/3 What is the slope of segment DA? D(4,-1) A(6,2)

OpenStudy (tylermckinney16):

Slope(m)= 1.5

Directrix (directrix):

Slope of CD =-2/3 Slope of DA = 3/2 Segment CD is perpendicular to Segment DA. There is a right angle at vertex D of the quadrilateral. ------------------ Segment AB has slope -2/3. Segment BC has slope 3/2. -2/3 * 3/2 = -1. If the product of the slopes of two lines is -1, then the lines are perpendicular. So, Segment AB is perpendicular to segment BC. There is a right angle at vertex B of the quadrilateral. ------- Segment DA is perpendicular to segment AB because the product of the slopes of the segments is -1. 3/2 * -2/3 = -1. There is a right angle at A in the quadrilateral. -------- Segment BC is perpendicular to segment CD because the product of the slopes is -1. 3/2 * -2/3 = -1 There is a right angle at vertex C of the quadrilateral. ------ The quadrilateral has 4 right angles which means that the quadrilateral is a rectangle.

OpenStudy (tylermckinney16):

That the answer?